OpenStudy (anonymous):

Harvey is 3 times as old as Jane. The sum of their ages is 48 years. Find the age of each.

OpenStudy (anonymous):

Harvey(H)= 3Jane(J) H+J=48 Substitute H=3J into the equation. 3J+J=48 Solve for J to find Harvey's age.

OpenStudy (anonymous):

Harvey's age = H Jane's age = J H = 3J (1) and J + H = 48 (2) plug (1) into (2) J + 3J = 48 4J = 48 J = 12 (3) plug (3) into (1) H =3 * 12 = 36